# Switti: Designing Scale-Wise Transformers for Text-to-Image Synthesis

### v3

@@ Abstract
Yandex Research                 22footnotemark: 2HSE University                 33footnotemark: 3MIPT                 44footnotemark: 4Skoltech                 55footnotemark: 5ITMO University
Figure 1: Switti produces high quality and aesthetic $512{\times}512$ image samples in around $0.13$ seconds.
@@ 1 Introduction
This work advances the family of scale-wise image generation models by introducing a novel transformer architecture for large-scale text-to-image generation. We begin by implementing next-scale prediction AR models for text-to-image generation, drawing on recent developments [73, 47]. However, the basic architecture encountered instabilities during training, resulting in suboptimal convergence. To address these issues, we introduce several architectural modifications to the transformer backbone, resulting in a strong next-scale prediction AR model for text-to-image synthesis.
Then, we investigate whether the scale-wise AR models require attending to all previous scales. We notice that an input image at the current resolution already contains information about all preceding scales by design. Therefore, we hypothesize that the model may not need explicit attention to these levels within its architecture. To test this, we remove the causal component from next-scale prediction transformers, enabling faster inference and improved scalability. In addition to the efficiency gains, we find that non-causal models deliver slightly better generative performance.
@@
  • We demonstrate that using a non-causal transformer makes Switti ${\sim}11\%$ more efficient for $512{\times}512$ image generation due to cheaper attention operations. Additionally, Switti reduces memory consumption during inference, previously needed for storing key-value (KV) cache, enabling better scaling to higher resolution image generation. Moreover, Switti slightly surpasses its causal counterpart in generation quality under the same training setups.
  • We find that Switti has weaker reliance on the text at high resolution scales. This observation allows us to disable classifier-free guidance at the last two steps, resulting in further ${\sim}20\%$ acceleration and better generation of fine-grained details, as confirmed by human evaluation.
@@ 3.1 Basic architecture
Our model adopts the pretrained RQ-VAE from VAR [73], which represents an image with $N{=}10$ scales. We slightly tune it on $512{\times}512$ resolution, as discussed in Section 4.3.
@@
The basic transformer architecture is adopted from VAR [73], where we incorporate cross-attention [76] layers between a self-attention layer and feed-forward network (FFN) in each transformer block. A pooled embedding from OpenCLIP ViT-bigG is propagated to the transformer blocks via Adaptive Layer Normalization (AdaLN) [82].
@@
Layer-normalization (LN) layers [1] are applied to the inputs of attention and FFN blocks. RMSNorm [90] layers are used for query-key (QK) normalization. We also use normalized 2D rotary positional embeddings (RoPE) [47, 22], which allow faster model adaptation to higher resolutions. The FFN block uses GeLU activation function [21]. The visualization of the basic architecture is in Appendix A.
@@
3.2 Training dynamics of the basic architecture
Here, we analyze the training performance of the basic model with $d{=}20$ transformer blocks and introduce the modifications improving the model stability and convergence.
We train the model in mixed-precision BF16/FP32 for $150$K iterations on the $256{\times}256$ image-text dataset described in Section 4.1. The detailed training and evaluation setups are in Appendix B. During the training, we track activation norms and standard metrics, such as FID [24], CLIPscore [23] and PickScore [34].
@@ 3.2 Training dynamics of the basic architecture
Figure 3: Last transformer block activation norms over training.
Figure 4: Evaluation metrics over training for various architectures.
To further reduce the growth of activation norms during training, we employ “sandwich”-like normalizations [12, 93], to keep the activation norms in a reasonable range. Specifically, we insert additional normalization layers right after each attention and feed-forward blocks and replace LN layers with RMSNorm for efficiency. As we show in Figure 3, this modification further mitigates the growth of activation norms during training, resulting in slightly better model performance, as evidenced in Figure 4 (Green).
Finally, following standard practices in language transformers [74, 13], we replace the GELU [21] activation in the FFN blocks with a SwiGLU [64] activation that allows the model to dynamically control the information flow via a learnable gating mechanism. Although SwiGLU is a prevalent choice in LLMs, we notice that it provides negligible effect on the final performance.
@@ 3.3 Employing a non-causal transformer
Next, we delve into the autoregressive next-scale prediction inference of the original VAR [73]. At each generation step $i$, the VAR transformer predicts a sequence of tokens $s_{i}$ conditioned on the previously generated tokens $s_{0},\dots,s_{i-1}$.The embeddings of these tokens form a feature map of the spatial size ${h_{i}{\times}w_{i}}$. This feature map is then upscaled to $h_{i+1}{\times}w_{i+1}$ and combined with previously predicted feature maps to serve as input for the next prediction step.
Therefore, we observe that the conditioning on the preceding scales occurs twice in VAR’s image generation: first, when forming the model’s input, and second, via the causal transformer. Based on this observation, we update the attention mask to allow self-attention layers to attend only to tokens at the current scale, as shown in Figure 5 (Right). This implies that the transformer is no longer causal, enabling more efficient sampling due to cheaper attention operations and eliminating the need for key-value (KV) cache. Figure 6 illustrates the next-scale prediction sampling using a non-causal transformer. Interestingly, this modification also slightly improves the performance in terms of CLIP score and PickScore, as shown in Figure 4.
Overall, we refer to the scale-wise model with the non-causal transformer architecture as Switti.
@@ 3.4 The role of text conditioning
Finally, we examine the effect of text conditioning at different model scales. Specifically, we analyze cross-attention maps and the model behavior when the textual prompt is switched during sampling.
@@
We plot a cross-attention map between image and text tokens, averaged across transformer blocks and image tokens at different scales. Figure 7 shows a typical pattern for a randomly selected prompt. The attention scores are primarily concentrated on the first and last tokens at most scales, while the highest scales tend to focus more on the beginning of the prompt. This pattern suggests that the model relies less on the prompt at higher scales. Such behavior is consistent across different prompts.
@@
Then, we investigate the impact of text-conditioning at various scales by switching the text prompt to a new one starting from a certain scale. The visualization of prompt switching is presented in Figure 8, with additional examples provided in Appendix C. Indeed, the prompt has minimal influence on the image semantics at the last two scales. Interestingly, switching a prompt at the middle scales results in a simple image blending approach.
Practical implications.
@@
Figure 7: Visualization of cross-attention maps at different scales for a random text prompt.
Figure 8: Textual prompt switching during image generation.
@@ 4.1 Pretraining
Data.
We collect the dataset of 100M image-text pairs that are prefiltered with a particular emphasis on relevance from the base set of 6B pairs from the web. We consider only images of sufficiently high aesthetic quality, based on the AADB [35] and TAD66k [20] aesthetic filters. We additionally consider sufficiently high-resolution images with at least 512px on each side. The dataset contains central crops of the images with an aspect ratio in $[0.75,1.33]$. The images are recaptioned using the LLaVA-v1.4-13B, LLaVA-v1.6-34B [45], and ShareGPT4V [9] models. The best caption is selected according to OpenCLIP ViT-G/14 [63].
Technical details.
Following the transformer scaling setup in VAR [73], we set the number of transformer layers $d{=}30$ for our main model, resulting in ${\sim}2.5$B trainable parameters.
During the first stage of pretraining, we train the model on $256{\times}256$ resolution using a batch size of $2,560$ for $400$K iterations that takes ${\sim}25$K NVIDIA A100 GPU hours. We start with a learning rate of $41e{-}$ 4 with a linear decay to $51e{-}$ 5. We use FSDP with a hybrid strategy for effective multi-host training and mixed precision BF16/FP32. To additionally reduce memory usage and speed up the training steps, we use precomputed textual embeddings from the text encoders.
Next, we train on $512{\times}512$ resolution for $200$K iterations using a batch size of $768$ and a learning rate of $51e{-}$ 5, linearly decaying to $75e{-}$ 7. This stage takes another ${\sim}12$K NVIDIA A100 GPU hours.
@@ 4.2 Supervised fine-tuning
The model is fine-tuned for $40$K iterations with a batch size of $80$ and a learning rate $75e{-}$ 7 on image central crops of resolution $512{\times}512$. We show the effect of supervised fine-tuning in Appendix D.
In addition, we slightly perturb the RQ-VAE latents prior the quantization step with Gaussian noise ($\sigma{=}0.01$) as an augmentation to mitigate overfitting. We observed that the perturbed latents after the quantization and dequantization steps produce almost identical images while resulting in ${\sim}70\%$ different tokens in a sequence.
@@ 4.3 RQ-VAE tuning
To compare the reconstruction quality of the original RQ-VAE and the one with a tuned decoder, we compute classic full-reference metrics PSNR, SSIM [80], LPIPS [92] and no-reference CLIP-IQA [77] metric on a held-out dataset of 300 images. Results presented in Table 1 demonstrate that the fine-tuned RQ-VAE outperforms an original model with respect to all metrics. Additionally, we provide several representative comparisons in Appendix E.
@@
Original 21.603 0.634 0.200 0.727
Fine-tuned 22.267 0.653 0.188 0.772
@@ 5 Experiments
SDXL [51] 0.867 2.6 0.226 0.360 0.77 14.4 0.217 0.384 0.78 7.6 0.55
SD3-medium [15] 0.934 2.0 0.227 0.354 1.01 19.5 0.215 0.363 0.91 13.1 0.65
Lumina-Next [93] 5.812 2.0 0.224 0.329 0.55 18.4 0.216 0.353 0.75 5.9 0.47
@@
LlamaGen [69] 3.821 0.8 0.208 0.274 -0.25 44.8 0.194 0.288 -0.45 26.9 0.32
HART [70] 0.063 0.7 0.223 0.341 0.75 20.9 0.216 0.366 0.84 5.8 0.55
Switti (AR) (ours) 0.139 2.5 0.227 0.354 0.91 17.8 0.217 0.378 0.87 9.6 0.61
Switti (ours) 0.127 2.5 0.227 0.356 0.95 17.6 0.217 0.381 0.91 9.5 0.63
@@
Figure 10: Qualitative comparison of Switti against the baselines.
@@ 5.1 Automated evaluation
To comprehensively evaluate the performance of our models, we use a combination of established automated metrics and a human preference study. For automated metrics, we report CLIPScore [23], ImageReward [83], PickScore [34], FID [24] and GenEval [18]. For all evaluated models, we generate images in their native resolution and resize them to $512{\times}512$. More details about the evaluation pipeline are in Appendix F.
@@ Results.
We present the metric values in Table 2. Switti achieves comparable performance to the baselines, ranking top-3 for 8 out of 9 automated metrics while exhibiting higher efficiency than most competitors. We do not provide automated metrics for Emu3 and Lumina-mGPT due to their exceptionally long sampling times, that makes generating 60,000 images infeasible in a reasonable time.
@@ 5.2 Human evaluation
While automated metrics are widely adopted for evaluating text-to-image models, we argue that they do not fully reflect all aspects of image generation quality. Therefore, we conduct a rigorous human preference study and consider it as our primary evaluation metric. This models are compared across several aspects: the presence of defects, textual alignment, image complexity and aesthetics quality. Our evaluators are specially trained experts who receive detailed and fine-grained instructions for each evaluation aspect. More details about the human evaluation setup are provided in Appendix G. For human evaluation, we generate two images for each of $128$ captions from Parti prompts collection [89], a set specifically curated for human preference study [15, 85, 61].
@@
For a side-by-side comparison, we generate images with classifier-free guidance set to 4 and deactivate it at the last two scales, as described in Section 3.4. We follow the original VAR inference implementation and apply Gumbel softmax sampling [28] with a decreasing temperature, starting from the third scale. At the first two scales, we use nucleus sampling with top-$k{=}400$ and top-$p{=}0.95$.
@@ Results.
We provide the results of a side-by-side comparison of Switti against the baselines for various image quality aspects in Figure 9. As follows from the human evaluation, Switti outperforms all AR baselines in most aspects, only lagging behind the much heavier models, Lumina-mGPT and Emu3, in terms of image complexity. As for diffusion models, DMD2 slightly outperforms Switti in terms of defects presence, which can be attributed to its adversarial training, whereas SD3 is better at text alignment, likely due to an additional text encoder. In other comparisons, Switti is on par with the diffusion models and their distilled versions, with respect to statistical error. We provide qualitative comparisons in Figures 10 and 15.
@@ 5.3 Inference performance evaluation
Next, we analyze the efficiency of Switti’s sampling and compare it to the baselines. We consider two settings: measurement of a single generator step without considering time for text encoding and VAE decoding; and inference time of a full text-to-image pipeline. All models are evaluated in half-precision with a batch size of $8$. KV-cache is enabled for all AR models. Since Switti’s generation resolution is currently limited to $512{\times}512$, we evaluate all models in this resolution to ensure a fair comparison. All models are evaluated on a single NVIDIA A100 80GB GPU.
@@
SDXL 2.6 25 12.4 0.867
SD3-medium 2.0 28 16.8 0.934
@@
Lumina-mGPT 7.0 1024 224.2
LlamaGen 0.8 1024 3.821
HART 0.7 10 4.7* 0.063
Switti (AR) 2.5 10 11.2* 0.139
Switti 2.5 10 9.5* 0.127
*Time averaged over 10 steps.
Table 3: Comparison of models’ $512{\times}512$ image generation time. All models are run in half-precision with a batch size of 8.
[…]

### v4

@@ Abstract
Yandex Research              22footnotemark: 2HSE University              33footnotemark: 3MIPT              44footnotemark: 4Skoltech              55footnotemark: 5AIRI              66footnotemark: 6ITMO
Figure 1: Switti produces high quality and aesthetic $1024{\times}1024$ image samples in around $0.5$ seconds.
@@ 1 Introduction
This work advances the family of scale-wise image generative models by introducing Switti, a next-scale prediction transformer for text-to-image generation, that rivals the performance of diffusion models, while being significantly faster. Drawing on recent developments [73, 46], we begin with implementing Switti (AR), exploring and mitigating stability issues in the training process.
Then, we investigate whether the scale-wise AR models require attending to all previous scales. We notice that an input image at the current resolution already contains information about preceding scales by design. Therefore, we hypothesize that the model may not need explicit attention to these levels within its architecture. To test this, we remove the causal component from next-scale prediction transformers, enabling faster inference and improved scalability. In addition to the efficiency gains, we find that non-causal models deliver slightly better generative performance.
@@
  • We demonstrate that using a non-causal transformer makes Switti ${\sim}21\%$ more efficient for $1024{\times}1024$ image generation due to cheaper attention operations. Additionally, Switti reduces memory consumption during inference, previously needed for storing key-value (KV) cache, by up to $2.3$ GB for a single image, enabling better scaling to higher resolution image generation. Moreover, Switti slightly surpasses its AR counterpart in generation quality under the same training setups.
  • We find that Switti has weaker reliance on the text at high resolution scales. This observation allows us to disable classifier-free guidance on the last two steps, resulting in further ${\sim}32\%$ acceleration and better generation of fine-grained details, as confirmed by human evaluation.
@@ 3.1 Architecture
Our model adopts the pretrained RQ-VAE from VAR [73], which represents an image with $N{=}10$ scales for resolutions $256$ and $512$, and $N{=}14$ for $1024$. We slightly tune RQ-VAE on $512{\times}512$ resolution, as discussed in Section 4.3.
@@
The transformer block architecture is adopted from VAR [73], where we incorporate cross-attention [76] layers between a self-attention layer and feed-forward network (FFN) in each transformer block. A pooled embedding from OpenCLIP ViT-bigG is propagated to the transformer blocks via Adaptive Layer Normalization (AdaLN) [82].
@@
RMSNorm [90] layers are applied to the inputs of attention and FFN blocks. Layer-normalization (LN) layers [1] are used for query-key (QK) normalization. We also use normalized 2D rotary positional embeddings (RoPE) [46, 21], which allow faster model adaptation to higher resolutions. We use SwiGLU activation function [63] in FFN blocks.
@@
3.2 Training stability issues
Here, we analyze the training performance of the basic model with $d{=}20$ transformer blocks and introduce a simple modification to improve training stability.
We train the model in mixed-precision BF16/FP32 for $150$K iterations on the $256{\times}256$ image-text dataset described in Section 4.1. The detailed training and evaluation setups are presented in Appendix A. During the training, we track activation norms and standard metrics, such as FID [23], CLIP Score [22] and PickScore [33].
@@ 3.2 Training stability issues
Figure 3: Last transformer block activation norms over training. Casting the prediction head to full-precision reduces the norm growth. “Sandwich”-normalization further mitigates the issue.
Figure 4: Evaluation of $d{=}20$ models on COCO 30K. Using the non-causal attention mask also slightly improves the performance.
To further reduce the growth of activation norms during training, we employ “sandwich”-like normalizations [12, 93], to keep the activation norms in a reasonable range. Specifically, we insert additional normalization layers right after each attention and feed-forward blocks. As we show in Figure 3, this modification further mitigates the growth of activation norms during training. However, we find that neither “sandwdich” normalizations nor the choice of normalization functions do not affect the metrics of the models at this scale, as evidenced in Appendix B.
@@ 3.3 Employing a non-causal transformer
Next, we delve into the autoregressive next-scale prediction sampling of the original VAR [73]. At each generation step $i$, the VAR transformer predicts a sequence of tokens $s_{i}$ conditioned on the previously generated tokens $s_{0},\dots,s_{i-1}$. The embeddings of these tokens form a feature map of the spatial size ${h_{i}{\times}w_{i}}$. This feature map is then upscaled to $h_{i+1}{\times}w_{i+1}$ and combined with the previously predicted feature maps to serve as an input for the next step.
Therefore, we observe that the conditioning on the preceding scales occurs twice in VAR’s image generation: first, implicitly, when forming the model input, and second, explicitly, using the causal transformer. Based on this observation, we update the attention mask to allow self-attention layers to attend only to tokens at the current scale, as shown in Figure 5 (Right). This implies that the transformer is no longer causal, enabling more efficient sampling due to cheaper attention operations and eliminating the need for key-value (KV) cache. Figure 6 illustrates the next-scale prediction sampling using a non-causal transformer. Overall, we refer to our scale-wise model with the non-causal transformer architecture as Switti.
A concurrent work, HART [69], makes a similar observation that self-attention in VAR is mostly local, allowing them to discard up to 80% tokens during training, while the model remains auto-regressive. In contrast, we remove causality in transformer and observe that this modification slightly improves performance in terms of CLIP Score and PickScore, as we show in Figure 4 (Green). Appendix C provides additional training loss analysis to justify our choice.
@@ 3.4 The role of text conditioning
Finally, we examine the effect of text conditioning at different model scales.
@@
First, we plot a cross-attention map between image tokens at different scales and text tokens (except the BOS-token), averaged across transformer blocks. We find that attention scores to the BOS-token are high for all scales and omit it to better visualize the pattern for other tokens. Figure 7 shows a typical pattern for two random prompts. At higher scales, the scores decrease for the EOS-token (used as a pooled embedding) but increase for the less semantic tokens, such as articles and stylistic words. This suggests a weaker reliance on the prompt at later scales. Such behavior is consistent across different prompts.
@@
Next, we investigate the impact of text-conditioning at various scales by switching the text prompt to a new one starting from a certain scale. The visualization of prompt switching is presented in Figure 8, with additional examples provided in Appendix D. Indeed, the prompt has minimal influence on the image semantics at the last two scales. Interestingly, switching a prompt at the middle scales results in a simple image blending approach.
Practical implication.
@@
Figure 7: Cross-attention maps between image tokens at each scale (Y-axis) and text tokens (X-axis). Attention scores decrease for main semantic tokens and pooled embeddings (EOS), while increase for less informative tokens, e.g., articles.
Figure 8: Prompt switching within $512{\times}512$ generation. Changing the prompt at late steps has a minor effect on the resulting image.
@@ 4.1 Pretraining
Following the transformer scaling setup in VAR [73], we set the number of transformer layers $d{=}30$ for our main model, resulting in ${\sim}2.5$B trainable parameters. We pretrain the model in two stages: first on images of $256{\times}256$ resolution, followed by training at $512{\times}512$.
As a training dataset, we collect $100$M image-text pairs that are filtered from a base set of $6$B pairs from the web. Images are filtered to ensure high aesthetic quality and recaptioned using open-source VLMs. We provide more technical details about data filtering and training hyperparameters in Appendix A.
@@ 4.2 Supervised fine-tuning
The model is fine-tuned for $40$K iterations with a batch size of $64$ and a learning rate $75e{-}$ 7 on image central crops of resolution $1024{\times}1024$.
In addition, we slightly perturb the RQ-VAE latents prior to the quantization step with Gaussian noise ($\sigma{=}0.01$) as an augmentation to mitigate overfitting. We observe that the perturbed latents after the quantization and dequantization steps produce visually indistinguishable images while resulting in ${\sim}80\%$ different tokens in a sequence, indicating that tokens at latter scales have less impact on the image quality.
@@ 4.3 RQ-VAE tuning
To compare the reconstruction quality of the original RQ-VAE and the one with a tuned decoder, we compute classic full-reference metrics PSNR, SSIM [80], LPIPS [92] and no-reference CLIP-IQA [77] metric on a held-out dataset of 300 images. Results in Table 1 demonstrate that the fine-tuned RQ-VAE outperforms an original model with respect to all metrics. Interestingly, the gains from fine-tuning RQ-VAE decoder for $512{\times}512$ resolution automatically transferred to $1024{\times}1024$ image reconstruction. Additionally, we provide several representative comparisons in Appendix E.
@@
$512{\times}512$ Reconstruction
Original 21.60 0.634 0.200 0.727
Fine-tuned 22.27 0.653 0.188 0.772
$1024{\times}1024$ Reconstruction
Original 22.53 0.703 0.207 0.683
Fine-tuned 23.91 0.733 0.177 0.748
@@ 5 Experiments
SDXL [50] 2.3 2.6 0.226 0.360 0.77 14.4 0.217 0.384 0.78 7.6 0.55
SD3-medium [15] 3.9 2.0 0.227 0.354 1.01 19.5 0.215 0.363 0.91 13.1 0.65
Lumina-Next [93] 5.8 2.0 0.224 0.329 0.55 18.4 0.216 0.353 0.75 5.9 0.47
@@
LlamaGen [68] 3.8 0.8 0.208 0.274 -0.25 44.8 0.194 0.288 -0.45 26.9 0.32
HART [70] 0.5 0.7 0.223 0.341 0.75 20.9 0.216 0.366 0.84 5.8 0.55
Switti 512 (ours) 0.1 2.5 0.227 0.356 0.95 17.6 0.217 0.381 0.91 9.5 0.62
Switti 1024 (ours) 0.5 2.5 0.229 0.355 0.96 18.2 0.219 0.378 0.81 8.1 0.62
@@
Figure 10: Qualitative comparison of Switti for $1024{\times}1024$ generation against the baselines. More examples are in Appendix H.
@@ 5.1 Automated evaluation
To comprehensively evaluate the performance of our models, we use a combination of established automated metrics and a human preference study. For automated metrics, we report CLIP Score [22], ImageReward [83], PickScore [33], FID [23] and GenEval [18]. For all evaluated models, we generate images in their native resolution and resize them to $512{\times}512$. More details about the evaluation pipeline are provided in Appendix F.
@@ Results.
We present the metric values in Table 2. Switti achieves comparable performance to the baselines, ranking top-3 for 7 out of 9 automated metrics while exhibiting higher efficiency than most competitors. We do not provide automated metrics for Emu3 and Lumina-mGPT due to their exceptionally long sampling times, that makes generating 60,000 images infeasible in a reasonable time.
@@ 5.2 Human evaluation
While automated metrics are widely adopted for evaluating text-to-image models, we argue that they do not fully reflect all aspects of image generation quality. Therefore, we conduct a rigorous human preference study across several aspects: the presence of defects, textual alignment, image complexity and aesthetics quality. Appendix G provides more details about the human evaluation setup. For human evaluation, we generate four images for each of $128$ captions from Parti prompts collection [89], a set specifically curated for human preference study [15, 85, 60].
@@
For a side-by-side comparison, we generate images with classifier-free guidance set to 6 and deactivate it at the last two scales, as described in Section 3.4. We follow the original VAR inference implementation and apply Gumbel softmax sampling [27] with a decreasing temperature, starting from the fifth scale. At the first four scales, we use nucleus sampling with top-$k{=}400$ and top-$p{=}0.95$.
@@ Results.
We provide the results of a side-by-side comparison of Switti against the baselines for various image quality aspects in Figure 9. As follows from the human evaluation, Switti outperforms all AR baselines in most aspects, only lagging behind Lumina-mGPT, Emu3 and LlamaGen in terms of image complexity. As for diffusion models, DMD2 slightly outperforms Switti in terms of defects presence, which can be attributed to its adversarial training, SD3-medium is better at text alignment, likely due to an additional text encoder, whereas SDXL surpasses our model in aesthetics. In other comparisons, Switti is on par with the diffusion models and their distilled versions, with respect to statistical error. We provide qualitative comparisons in Figures 10 and 15.
@@ 5.3 Inference performance evaluation
Next, we analyze the efficiency of Switti’s sampling and compare it to the baselines. We consider two settings: measurement of a single generator step without considering time for text encoding and VAE decoding; and inference time of a full text-to-image pipeline. All models are evaluated in half-precision with a batch size of $8$. KV-cache is enabled for all AR models. All models are evaluated on a single NVIDIA A100 80GB GPU.
@@
SDXL-Turbo 2.6 4 42.1 0.4
DMD2 2.6 4 42.1 0.4
SDXL 2.6 25 42.1 2.3
SD3-medium 2.0 28 61.3 3.9
@@
Lumina-mGPT 7.0 1024 527.2
LlamaGen 0.8 1024 3.8*
HART 0.7 14 4.9** 0.5
Switti (AR) 2.5 14 33.2** 0.6
Switti 2.5 14 26.4** 0.5
*Can only generate in 512x512.
**Time averaged over 14 steps.
Table 3: Comparison of $1024{\times}1024$ image generation times.
As follows from Table 3, Switti is among the most efficient image generation models of similar size, being more than 4 times faster than SDXL. Notably, Switti takes the same time as HART to generate a batch of $1024{\times}1024$ images while being more than three times larger. This efficiency stems from the fact that we do not employ an additional diffusion model during de-tokenization in RQ-VAE, from the transitioning to a non-causal architecture, and from disabling classifier-free guidance at the latest scales. Table 8 presents the evaluation on $512{\times}512$ resolution.